# Berry phases and connection matrices defined on homogeneous spaces attached to Jacobi groups

### v2

@@ 1. Introduction
We recall here our interrest to find a geometric significance to the phase of the scalar product of coherent states [79, 71, 75]. The answer to this question was given by Pancharatnam for the Poincaré sphere [78, 84], see also [1] and [74, Proposition 5.1] in the language of holonomy (see § 5.1.3) of a loop in the projective Hilbert space, and by Perelomov [79, page 63] for the sphere $\text{S}^{2}=\frac{\text{SU}(2)}{\text{U}(1)}$. A general answer to this question using the coherent state embedding and the so called ” Cauchy formulas” was given in [11] and [37]. We also studied this problem in [9][12]. Explicit calculation was presented for the compact Grassmann manifold $G_{n}(\mathbb{C}^{m+n})=\frac{\text{SU}(n+m)}{\text{S}(\text{U}(n)\times\text{U}(m))}$ in [11], where it was proved that the phase of the scalar product of two coherent states is twice the symplectic area of a geodesic triangle determined by the corresponding points on the manifold and the origin of the system of coordinates, see also [51, Theorem 2.1]. The same result is also true for the noncompact dual $\frac{\text{SU}(n,m)}{\text{S}(\text{U}(n)\times\text{U}(m))}$ of the compact Grassmann manifold [9, 10]. In [21] the change of coordinates $x\rightarrow z$ in (3.1) below was called $FC$-transform (fundamental conjecture [86, 52, 55]). We observed that [31, Remark 3]
@@
Proposition 1 and Comment 1 are improved versions of older results, while Remark 3 compare our approach to Berry phase on Kähler  manifolds [35, 31, 21] with the geometric phase in [84, 83, 50]. The new relevant results presented in this paper are § 3.4, formula (3.47) of $\theta_{{\mathcal{X}}^{J}_{1}}(x,y,q,p)$, formula (3.66) of Berry phase on ${\mathcal{X}}^{J}_{1}$ in $(u,v)=(m+\operatorname{i}n,x+\operatorname{i}y)$, formulae (3.67), (3.68), (3.69) of the Berry phase in $(w,v)=(\alpha+\operatorname{i}\beta,x+\operatorname{i}y)$, (3.70) for the Berry phase in $(x,y,q,p)$, Lemma 2 which gives $\omega_{{\mathcal{X}}^{J}_{n}}(x,y,p,q)$, formula (4.20) $\theta^{\prime}_{\tilde{{\mathcal{X}}}^{J}_{1}}(x,y,q,p,\kappa)$ of the connection matrix on $\tilde{{\mathcal{X}}^{J}_{1}}$, the covariant derivatives $Dx$,( $Dy,~{}Dq,~{}Dp$) (4.25), ( (4.26), (4.27), respectively (4.28)) on ${\mathcal{X}}^{J}_{1}$, formulae of $Dx$, ($Dy,~{}Dq,~{}Dp,~{}D\kappa$) (4.29) ((4.30), (4.31), (4.32), respectively (4.33)) on $\tilde{{\mathcal{X}}}^{J}_{1}$.
@@
Notation We denote by $\mathbb{R}$, $\mathbb{C}$, $\mathbb{Z}$ and $\mathbb{N}$ the field of real numbers, the field of complex numbers, the ring of integers, and the set of non-negative integers, respectively. We denote the imaginary unit $\sqrt{-1}$ by $\operatorname{i}$, the real and imaginary parts of a complex number $z\in\mathbb{C}$ by $\operatorname{Re}z$ and $\operatorname{Im}z$ respectively, and the complex conjugate of $z$ by $\bar{z}$. We denote by ${\operatorname{d}}$ the differential. We use Einstein’s summation convention, i.e. repeated indices are implicitly summed over. The set of vector fields (1-forms) on real manifolds is denoted by ${\mathfrak{D}}^{1}$ (respectively ${\mathfrak{D}}_{1}$). We denote a mixed tensor contravariant of degree $r$ and covariant of degree $s$ by ${\mathfrak{D}}^{r}_{s}={\mathfrak{D}}^{r}\times{\mathfrak{D}}_{s}$, where ${\mathfrak{D}}^{r}=\underbrace{{\mathfrak{D}}^{1}\times\dots\times{\mathfrak{D}}^{1}}_{r}$ and ${\mathfrak{D}}_{s}=\underbrace{{\mathfrak{D}}_{1}\times\dots\times{\mathfrak{D}}_{1}}_{s}$ [60, pages 13-17]. If $M$ is a complex manifold we denote by ${\mathfrak{A}}^{r,s}$ the tensor fields of type $(r,s)$. If we denote with Roman capital letteres the Lie groups, then their associated Lie algebras are denoted with the corresponding lower-case letteres. If ${\mathfrak{H}}$ is a Hilbert space, than we adopt the convention that the scalar product (,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) on ${\mathfrak{H}}\times{\mathfrak{H}}$ is antilinear in the first factor $(\lambda a,b)=\bar{\lambda}(a,b),\quad\lambda\in\mathbb{C}\setminus\{0\}$. If $\pi$ is a representation of a Lie grup $G$ on the Hilbert ${\mathfrak{H}}$ and $X\in{\mathfrak{g}}$, then we denote $\bf{X}:=\operatorname{d}\pi(X)$ [14, 22, 79]. The interior product $i_{X}\omega$ (interior multiplication or contraction) of the differential form $\omega$ with $X\in{\mathfrak{D}}^{1}$ is denoted $X\lrcorner\omega$. We denote by $M(n,m,{\mathbb{F}})$ the set of $n\times m$ matrices with elements in the field ${\mathbb{F}}$ and $M(n,{\mathbb{F}})$ denotes $M(n,n,{\mathbb{F}})$. If $X\in M(n,m,{\mathbb{F}})$, then $X^{t}$ denotes the transpose of $X$. We denote by $MS(n,{\mathbb{F}})=\{X\in M(n,{\mathbb{F}})|X=X^{t}\}$. The conjugate transpose (or hermitian transpose) of $A\in M(q,\mathbb{C})$ is $A^{H}:=\bar{A}^{t}$, also denoted $A^{*}$, $A^{\dagger}$, $A^{+}$. If $f$ is a function on $\mathbb{C}^{n}$, we write for the total differential of $f$ $\operatorname{d}f=\partial f+\bar{\partial}f$, $\partial f=\sum_{1}^{n}{\partial_{\alpha}f}\operatorname{d}z_{\alpha}$, where $\partial_{\alpha}f=\frac{\partial f}{\partial z_{\alpha}}$ [59, page 6]. If $f$ is a complex function, then by $f-{cc}$ we mean $f-\bar{f}$.
@@ 2. Preparation
(2.9) $e_{z,w}:=e^{\sqrt{\mu}z{{\mbox{\boldmath{$a$}}}}^{\dagger}+w{{\mbox{\boldmath{$K$}}}}_{+}}e_{0},~{}z\in\mathbb{C},~{}|w|<1,$
@@
(2.11) $-\operatorname{i}\omega_{{\mathcal{D}}^{J}_{1}}(w,z)\!=\!\frac{2k}{P^{2}}\operatorname{d}w\wedge\operatorname{d}\bar{w}\!+\!\nu\frac{{\mathcal{A}}\wedge\bar{{\mathcal{A}}}}{P},P:=1-|w|^{2},{\mathcal{A}}={\mathcal{A}}(w,z):=\operatorname{d}z\!+\!\bar{\eta}\operatorname{d}w.$
@@
(2.19) $-\operatorname{i}\omega(v,\eta)=\frac{k}{2y^{2}}\operatorname{d}v\wedge\operatorname{d}\bar{v}+\nu\{-2\operatorname{i}[\frac{\bar{\eta}}{(v+\operatorname{i})^{2}}\operatorname{d}v\wedge\operatorname{d}\bar{\eta}+\frac{\eta}{(\bar{v}-\operatorname{i})^{2}}\operatorname{d}\bar{v}\wedge\operatorname{d}\eta]+\operatorname{d}\eta\wedge\operatorname{d}\bar{\eta}\},$
@@
The Berndt–Kähler’s two-form (symplectic two-form) invariant to the action of $G^{J}(\mathbb{R})_{0}$ $=\rm{SL}(2,\mathbb{R})\ltimes\mathbb{C}$, is (2.31a) ((2.31c))
@@
[…]
@@
$-\operatorname{i}\omega_{{\mathcal{X}}^{J}_{1}}(\tau,z)=f^{\prime}_{\tau\bar{\tau}}\operatorname{d}\tau\wedge\operatorname{d}\bar{\tau}+f^{\prime}_{\tau\bar{z}}\operatorname{d}\tau\wedge\operatorname{d}\bar{z}-\bar{f}^{\prime}_{\tau\bar{z}}\operatorname{d}\bar{\tau}\wedge\operatorname{d}z+f^{\prime}_{z\bar{z}}\operatorname{d}z\wedge\operatorname{d}\bar{z}.$
@@
$-\operatorname{i}\omega(v,\eta)=h_{v\bar{v}}\operatorname{d}v\wedge\operatorname{d}\bar{v}+h_{v\bar{\eta}}\operatorname{d}v\wedge\operatorname{d}\bar{\eta}-\bar{h}_{v\bar{\eta}}\operatorname{d}\bar{v}\wedge\operatorname{d}\eta+h_{\eta\bar{\eta}}\operatorname{d}\eta\wedge\operatorname{d}\bar{\eta}.$
@@
(2.55) $-\operatorname{i}\omega_{{\mathcal{X}}^{J}_{n}}(v,u)=\frac{k}{2}\operatorname{tr}(H\wedge\bar{H})+\frac{2\nu}{\operatorname{i}}\operatorname{tr}(G^{t}D\wedge\bar{G}),\quad D:=(\bar{v}-v)^{-1},~{}H:=D\operatorname{d}v.$
@@ 3.1. Balanced metric
(3.1) $\underline{e}_{x}=\exp(\sum_{\phi\in\Delta^{+}}x_{\phi}{{\mbox{\boldmath{$X$}}}}^{+}_{\phi}-{\bar{x}}_{\phi}{{\mbox{\boldmath{$X$}}}}^{-}_{\phi})e_{0},\quad e_{z}=\exp(\sum_{\phi\in\Delta^{+}}z_{\phi}{{\mbox{\boldmath{$X$}}}}^{+}_{\phi})e_{0},$
@@
(3.2a) $-\operatorname{i}\omega_{M}$ =¯$=\partial\bar{\partial}f,~{}f(z,\bar{z})=\log K(z,\bar{z}),~{}K(z,\bar{z}):=(e_{{z}},e_{{z}}),$
@@
(3.6) $-\operatorname{i}\omega_{M}=\frac{\operatorname{i}}{2}\sum_{\alpha,\beta}h_{\alpha\bar{\beta}}\operatorname{d}z_{\alpha}\wedge\operatorname{d}\bar{z}_{\beta}.$
@@ 3.2. Berry phase on homogenous Kähler  manifolds
3.3. Linear Hamiltonian in the generators of the Jacobi group $G^{J}_{1}(\mathbb{R})$
@@
Let us consider a linear Hermitian Hamiltonian $H$ in the generators of the Jacobi group $G^{J}_{1}$
@@
[…]
@@
(3.26a) $\frac{\partial H(x,y)}{\partial x}=$ $\frac{2k}{y}[(m\!+\!c)x\!+\!n];\frac{\partial H(x,y)}{\partial y}=\!\frac{k}{y^{2}}[(m\!+\!c)(y^{2}-x^{2})\!+\!m\!-\!c\!-\!2nx];$
(3.26b) $\frac{\partial H(q,p)}{\partial q}=$ $2\nu[(m\!+\!c)q\!-\!np\!+\!a];\frac{\partial H(q,p)}{\partial p}\!=\!2\nu[(c\!-\!m)p\!-\!nq\!+\!b].$
@@
(3.27a) $\dot{q}_{1}$ $=k\dot{x}=A_{1}=y^{2}\frac{\partial H}{\partial y}=k[(m+c)(y^{2}-x^{2})+m-c-2nx]$
(3.27b) $\dot{q}_{2}$ $=2\nu\dot{q}=A_{2}=\frac{\partial H}{\partial p}-\frac{q}{2\nu}\frac{\partial H}{\partial\kappa}=2\nu[{(c-m)p-nq+b}-q]-q\frac{\partial H}{\partial\kappa}$
(3.27c) $\dot{p}_{1}$ $=y^{-2}\dot{y}=B_{1}=-\frac{1}{k}\frac{\partial H}{\partial x}=-\frac{1}{y}[(c+m)x+n]$
@@
(3.28) $-\operatorname{i}\omega(w,\bar{w})=\frac{2k}{P^{2}}\operatorname{d}w\wedge\operatorname{d}\bar{w}.$

### v3

@@ 1. Introduction
We recall here our interrest to find a geometric significance to the phase of the scalar product of coherent states [79, 71, 75]. The answer to this question was given by Pancharatnam for the Poincaré sphere [78, 84], see also [1] and [74, Proposition 5.1] in the language of holonomy (see § 5.1.3) of a loop in the projective Hilbert space, and by Perelomov [79, page 63] for the sphere $\text{S}^{2}=\frac{\text{SU}(2)}{\text{U}(1)}$. A general answer to this question using the coherent state embedding and the so called ” Cauchy formulas” was given in [11] and [37]. We also studied this problem in [9][12]. Explicit calculation was presented for the compact Grassmann manifold $G_{n}(\mathbb{C}^{m+n})=\frac{\text{SU}(n+m)}{\text{S}(\text{U}(n)\times\text{U}(m))}$ in [11], where it was proved that the phase of the scalar product of two coherent states is twice the symplectic area of a geodesic triangle determined by the corresponding points on the manifold and the origin of the system of coordinates, see also [51, Theorem 2.1]. The same result is also true for the noncompact dual $\frac{\text{SU}(n,m)}{\text{S}(\text{U}(n)\times\text{U}(m))}$ of the compact Grassmann manifold [9, 10]. In [21] the change of coordinates $x\rightarrow z$ in (3.1) below was called FCFC-transform (fundamental conjecture [86, 52, 55]). We observed that [31, Remark 3]
@@
Proposition 1 and Comment 1 are improved versions of older results, while Remark 3 compare our approach to Berry phase on Kähler  manifolds [35, 31, 21] with the geometric phase in [84, 83, 50]. The new relevant results presented in this paper are § 3.4, formula (3.47) of $\theta_{{\mathcal{X}}^{J}_{1}}(x,y,q,p)$, formula (3.66) of Berry phase on ${\mathcal{X}}^{J}_{1}$ in $(u,v)=(m+\operatorname{i}n,x+\operatorname{i}y)$, formulae (3.67), (3.68), (3.69) of the Berry phase in $(w,v)=(\alpha+\operatorname{i}\beta,x+\operatorname{i}y)$, (3.70) for the Berry phase in $(x,y,q,p)$, Lemma 2 which gives $\omega_{{\mathcal{X}}^{J}_{n}}(x,y,p,q)$, formula (4.20) $\theta^{\prime}_{\tilde{{\mathcal{X}}}^{J}_{1}}(x,y,q,p,\kappa)$ of the connection matrix on $\tilde{{\mathcal{X}}^{J}_{1}}$, the covariant derivatives DxDx,( Dy,Dq,DpDy,~Dq,~Dp) (4.25), ( (4.26), (4.27), respectively (4.28)) on ${\mathcal{X}}^{J}_{1}$, formulae of DxDx, ($Dy,~Dq,~Dp,~D\kappa$) (4.29) ((4.30), (4.31), (4.32), respectively (4.33)) on $\tilde{{\mathcal{X}}}^{J}_{1}$.
@@
Notation We denote by $\mathbb{R}$, $\mathbb{C}$, $\mathbb{Z}$ and $\mathbb{N}$ the field of real numbers, the field of complex numbers, the ring of integers, and the set of non-negative integers, respectively. We denote the imaginary unit $\sqrt{-1}$ by $\operatorname{i}$, the real and imaginary parts of a complex number $z\in\mathbb{C}$ by $\operatorname{Re}z$ and $\operatorname{Im}z$ respectively, and the complex conjugate of $z$ by $\bar{z}$. We denote by $\operatorname{d}$} the differential. We use Einstein’s summation convention, i.e. repeated indices are implicitly summed over. The set of vector fields (1-forms) on real manifolds is denoted by ${\mathfrak{D}}^{1}$ (respectively ${\mathfrak{D}}_{1}$). We denote a mixed tensor contravariant of degree $r$ and covariant of degree $s$ by ${\mathfrak{D}}^{r}_{s}={\mathfrak{D}}^{r}\times{\mathfrak{D}}_{s}$, where ${\mathfrak{D}}^{r}=\underbrace{{\mathfrak{D}}^{1}\times\dots\times{\mathfrak{D}}^{1}}_{r}$ and ${\mathfrak{D}}_{s}=\underbrace{{\mathfrak{D}}_{1}\times\dots\times{\mathfrak{D}}_{1}}_{s}$ [60, pages 13-17]. If $M$ is a complex manifold we denote by ${\mathfrak{A}}^{r,s}$ the tensor fields of type $(r,s)$. If we denote with Roman capital letteres the Lie groups, then their associated Lie algebras are denoted with the corresponding lower-case letteres. If $\mathfrak{H}$} is a Hilbert space, than we adopt the convention that the scalar product $(\cdot,\cdot)$ on $\mathfrak{H}$$\times{\mathfrak{H}}$ is antilinear in the first factor $(\lambda a,b)=\bar{\lambda}(a,b),\quad\lambda\in\mathbb{C}\setminus\{0\}$. If $\pi$ is a representation of a Lie grup $G$ on the Hilbert $\mathfrak{H}$} and $X\in{\mathfrak{g}}$, then we denote $\bf{X}:=\operatorname{d}\pi(X)$ [14, 22, 79]. The interior product $i_{X}\omega$ (interior multiplication or contraction) of the differential form $\omega$ with $X\in{\mathfrak{D}}^{1}$ is denoted $X\lrcorner\omega$. We denote by $M(n,m,{\mathbb{F}})$ the set of $n\times m$ matrices with elements in the field $\mathbb{F}$} and $M(n,{\mathbb{F}})$ denotes $M(n,n,{\mathbb{F}})$. If $X\in M(n,m,{\mathbb{F}})$, then $X^{t}$ denotes the transpose of $X$. We denote by $MS(n,{\mathbb{F}})=\{X\in M(n,{\mathbb{F}})|X=X^{t}\}$. The conjugate transpose (or hermitian transpose) of $A\in M(q,\mathbb{C})$ is $A^{H}:=\bar{A}^{t}$, also denoted $^{*}, AA^{\dagger}$, $A^{+}$. If $f$ is a function on $\mathbb{C}^{n}$, we write for the total differential of $f$ $\operatorname{d}f=\partial f+\bar{\partial}f$, $\partial f=\sum_{1}^{n}{\partial_{\alpha}f}\operatorname{d}z_{\alpha}$, where $\partial_{\alpha}f=\frac{\partial f}{\partial z_{\alpha}}$ [59, page 6]. If $f$ is a complex function, then by fccf-{cc} we mean $f-\bar{f}$.
@@ 2. Preparation
(2.9) $e_{z,w}:=e^{\sqrt{\mu}z{{\mbox{{$a$}}}}^{\dagger}+w{{\mbox{{$K$}}}}_{+}}e_{0},~z\in\mathbb{C},~|w|<1,$
@@
(2.11) $\operatorname{i}\omega_{{\mathcal{D}}^{J}_{1}}(w,z)\!=\!\frac{2k}{P^{2}}\operatorname{d}w\wedge\operatorname{d}\bar{w}\!+\!\nu\frac{{\mathcal{A}}\wedge\bar{{\mathcal{A}}}}{P},P:=1-|w|^{2},{\mathcal{A}}={\mathcal{A}}(w,z):=\operatorname{d}z\!+\!\bar{\eta}\operatorname{d}w.$
@@
(2.19) $\operatorname{i}\omega(v,\eta)=\frac{k}{2y^{2}}\operatorname{d}v\wedge\operatorname{d}\bar{v}+\nu\{-2\operatorname{i}[\frac{\bar{\eta}}{(v+\operatorname{i})^{2}}\operatorname{d}v\wedge\operatorname{d}\bar{\eta}+\frac{\eta}{(\bar{v}-\operatorname{i})^{2}}\operatorname{d}\bar{v}\wedge\operatorname{d}\eta]+\operatorname{d}\eta\wedge\operatorname{d}\bar{\eta}\},$
@@
The Berndt–Kähler’s two-form (symplectic two-form) invariant to the action of $G^{J}(\mathbb{R})_{0}$ $\rm{SL}(2,\mathbb{R})\ltimes\mathbb{C}$, is (2.31a) ((2.31c))
@@
(2.34) ${\rm FC}_{1}\colon\ 2\operatorname{i}u=(v+\operatorname{i})\eta-(v-\operatorname{i})\bar{\eta},\qquad{\rm FC}^{-1}_{1}\colon\eta=\frac{u\bar{v}-\bar{u}v}{\bar{v}-v}+\operatorname{i}r,$
@@
$\operatorname{i}\omega_{{\mathcal{X}}^{J}_{1}}(\tau,z)=f^{\prime}_{\tau\bar{\tau}}\operatorname{d}\tau\wedge\operatorname{d}\bar{\tau}+f^{\prime}_{\tau\bar{z}}\operatorname{d}\tau\wedge\operatorname{d}\bar{z}-\bar{f}^{\prime}_{\tau\bar{z}}\operatorname{d}\bar{\tau}\wedge\operatorname{d}z+f^{\prime}_{z\bar{z}}\operatorname{d}z\wedge\operatorname{d}\bar{z}.$
@@
$\operatorname{i}\omega(v,\eta)=h_{v\bar{v}}\operatorname{d}v\wedge\operatorname{d}\bar{v}+h_{v\bar{\eta}}\operatorname{d}v\wedge\operatorname{d}\bar{\eta}-\bar{h}_{v\bar{\eta}}\operatorname{d}\bar{v}\wedge\operatorname{d}\eta+h_{\eta\bar{\eta}}\operatorname{d}\eta\wedge\operatorname{d}\bar{\eta}.$
@@
(2.55) $\operatorname{i}\omega_{{\mathcal{X}}^{J}_{n}}(v,u)=\frac{k}{2}\operatorname{tr}(H\wedge\bar{H})+\frac{2\nu}{\operatorname{i}}\operatorname{tr}(G^{t}D\wedge\bar{G}),\quad D:=(\bar{v}-v)^{-1},~H:=D\operatorname{d}v.$
@@ 3.1. Balanced metric
(3.1) $\underline{e}_{x}=\exp(\sum_{\phi\in\Delta^{+}}x_{\phi}{{\mbox{{$X$}}}}^{+}_{\phi}-{\bar{x}}_{\phi}{{\mbox{{$X$}}}}^{-}_{\phi})e_{0},\quad e_{z}=\exp(\sum_{\phi\in\Delta^{+}}z_{\phi}{{\mbox{{$X$}}}}^{+}_{\phi})e_{0},$
@@
(3.2a) $-\operatorname{i}\omega_{M}$ $=\partial\bar{\partial}f,~f(z,\bar{z})=\log K(z,\bar{z}),~K(z,\bar{z}):=(e_{{z}},e_{{z}}),$
@@
(3.6) $\operatorname{i}\omega_{M}=\frac{\operatorname{i}}{2}\sum_{\alpha,\beta}h_{\alpha\bar{\beta}}\operatorname{d}z_{\alpha}\wedge\operatorname{d}\bar{z}_{\beta}.$
@@ 3.2. Berry phase on homogenous Kähler  manifolds
$3.$ Linear Hamiltonian in the generators of the Jacobi group $G^{J}_{1}(\mathbb{R})$
@@
(3.19) ${\mbox{{$H$}}}=\epsilon_{a}{\mbox{{$a$}}}+\bar{\epsilon}_{a}{\mbox{{$a$}}}^{\dagger}+\epsilon_{0}{\mbox{{$K$}}}_{0}+\epsilon_{+}{\mbox{{$K$}}}_{+}+\epsilon_{-}{\mbox{{$K$}}}_{-},$
@@
(3.24a) ${\mbox{{$a$}}}=\frac{\partial}{\sqrt{\mu}\partial z};~{\mbox{{$a$}}}^{\dagger}=\sqrt{\mu}z+w\frac{\partial}{\sqrt{\mu}\partial z};$
@@
(3.26a) $\frac{\partial H(x,y)}{\partial x}=$ $\frac{2k}{y}[(m\!+\!c)x\!-\!n];\frac{\partial H(x,y)}{\partial y}=\!\frac{k}{y^{2}}[(m\!+\!c)(y^{2}-x^{2})\!+\!m\!-\!c\!+\!2nx];$
(3.26b) $\frac{\partial H(q,p)}{\partial q}=$ $2\nu[(m\!+\!c)q\!+\!np\!+\!a];\frac{\partial H(q,p)}{\partial p}\!=\!2\nu[(c\!-\!m)p\!+\!nq\!+\!b].$
@@
(3.27a) $\dot{q}_{1}$ $=k\dot{x}=A_{1}=y^{2}\frac{\partial H}{\partial y}=k[(m+c)(y^{2}-x^{2})+m-c+2nx]$
(3.27b) $\dot{q}_{2}$ $=2\nu\dot{q}=A_{2}=\frac{\partial H}{\partial p}-\frac{q}{2\nu}\frac{\partial H}{\partial\kappa}=2\nu[{(c-m)p-nq+b}-q]-\frac{q}{2\nu}\frac{\partial H}{\partial\kappa}$
(3.27c) $\dot{p}_{1}$ $=y^{-2}\dot{y}=B_{1}=-\frac{1}{k}\frac{\partial H}{\partial x}=-\frac{1}{y}[(c+m)x-n]$
@@
(3.28) $\operatorname{i}\omega(w,\bar{w})=\frac{2k}{P^{2}}\operatorname{d}w\wedge\operatorname{d}\bar{w}.$